# Limits on Dark Matter Compact Objects implied by Supermagnified Stars in Lensing Clusters

### v1

@@ 4 Results
1 a MACS J1149 Icarus (Kelly et al., 2018) 0.54 1.49 0.832 12.46 0.0080 0.13 25-26 (F125W)
2 b MACS J0416 Spock-NW (Rodney et al., 2018) 0.397 1.0054 0.742 10.29 0.0028 25-26.5 (F814W))
@@
4 c MACS J0416 Warhol (Kaurov et al., 2019) 0.397 0.94 0.686 8.74 0.02 0.06 26.25 (F125W)
5 d WHL0137–08 Earendel (Welch et al., 2022a) 0.566 6.2 0.590 24.01 0.0060 0.009 27.2 (F150W)
6 e PSZ1 G311.65 Godzilla (Diego et al., 2022) 0.443 2.37 0.55 $\approx$ 22 (F814W)
7 f Abell 2744 LS7 (Chen et al., 2022) 0.308 2.65 0.726 15.68 0.0052 0.15 27.05 (F150W)
8 g MACS J0647 star-1 (Meena et al., 2023b) 0.591 4.8 0.544 29.28 0.1 27.343 (F227W)
9 g MACS J0647 star-2 (Meena et al., 2023b) 0.591 4.8 0.549 30.59 0.32 28.330 (F227W)
10 h El Gordo Quyllur (Diego et al., 2023b) 0.870 2.1878 0.662 23.84 7.5 $\cdot 10^{-3}$ 25.5 (F356W)
11 i Abell 370 LS11 (Meena et al., 2023a) 0.375 1.2567 0.703 18.89 0.0017 0.6 29.51 (F200LP)
12 j MACS J0416 Mothra (Diego et al., 2023a) 0.397 2.091 0.936 15.78 0.05 $\approx$ 27.8 (F200W)
Table 1: List of 12 supermagnified star candidates, with lens and source redshifts $z_{l}$ and $z_{s}$, convergence at image position $\kappa_{0}$, eigenvalue inverse gradient $\mathbf{g}=\mathbf{\nabla}(-\kappa-\gamma$) at the macro critical curve (MaCL), angular separation from source to MaCL $r$, and observed AB magnitude of each transient at given filter (AB mag). The values of $\kappa_{0}$ and $g^{-1}$ are the average of the set of publicly available lens models for each lens.
@@
    Caminha (v4) (Caminha et al. (2017a)), GLAFIC (v4) (Oguri (2010), Kawamata et al. (2016), Kawamata et al. (2018)), Keeton (v4) (McCully et al. (2014), Ammons et al. (2014), Keeton (2010)), Sharon (v4 Cor.) (Johnson et al. (2014), Jullo et al. (2007)), Zitrin-ltm-gauss (v3) and Zitrin-nfw (v3) (Zitrin et al. (2009), Zitrin et al. (2013)).
@@
We now proceed to a maximum likelihood estimate of the optimal value of $\kappa_{\star}$ to explain the distribution of separations in Figure 1, and an upper limit to the fraction of dark matter that may be composed of compact objects. Initially, we assume for simplicity that any compact objects accounting for part of the dark matter have a similar mass distribution as the intracluster stars, and in the next section we discuss the range of masses that our limit actually applies to.
@@
where $p_{i}(r_{i},g_{i},\kappa_{\text{b}})$ is the probability density of observing a stellar microlensing event at a distance $r_{i}$ from the MaCL of event $i$, given a microlensing band of width $r_{w,i}=(\kappa_{\star,i}+\kappa_{\text{b}})/g_{i}$. We are assuming that the total contribution from point masses to microlensing is the sum of the known intracluster light ($\kappa_{\star}$) and a hypothetical contribution from dark matter compact objects, such as primordial black holes ($\kappa_{\text{b}}$). We then find the optimal value of $\kappa_{\text{b}}$ that maximizes this likelihood function.
@@
The result for the likelihood function is shown in Figure 2. The optimal value of the dark matter surface density is $\kappa_{b}\simeq 0.95\times 10^{-3}$, and is consistent with zero. Therefore, the observed distribution of supermagnified star images is consistent with no compact objects in the dark matter, with the known stellar population in the intracluster light being ideal for explaining the observations. Figure 2 shows also the $(1,2,3)\sigma$ upper limits that can be placed on any additional compact objects, if we treat the $\ln\mathcal{L}$ function as a $\chi^{2}$ function. The $3\sigma$ upper limit, at a confidence level of 0.3%, is $\kappa_{b}<0.019$. We warn that this upper limit includes only the statistical uncertainty associated with the 9 events we used, but we have not attempted to incorporate other uncertainties due to the modeling of the lenses or the mass function of the stars in the intracluster light. These values for the upper limits to the fraction of dark matter in compact objects should also be slightly increased when we take into account that $\sim$ 10% of the total mass contributing to lensing is in the hot, X-ray emitting gas (note that the gas usually has a density profile that is less centrally concentrated than the dark matter in the strong lensing region).
@@
As discussed at the beginning of Section 3, we have excluded from this analysis event 6, which similarly to event 11 has a magnified image further from the model MaCL by a factor $\sim 4$ than the predicted width $r_{w}$ for $\kappa_{\star}=0.01\kappa_{0}$, because of the model uncertainty in the true separation from the MaCL. Had we included this event, the $3\sigma$ upper limit on $\kappa_{b}$ increases only to $\kappa_{b}<0.022$.
@@
As a consequence, we can establish an upper limit to the fraction of dark matter that may be in compact objects. The average local convergence in the 9 events we have used is $\bar{\kappa}_{0}=0.691$, so the $3\sigma$ upper limit on the dark matter fraction in compact objects is $\kappa_{b}/\bar{\kappa}_{0}\simeq 2.8\%$, which is only slightly increased to $\simeq 3\%$ when taking into account the contribution of hot gas to $\bar{\kappa}_{0}$.
@@
Figure 2: Likelihood function $\ln\mathcal{L}$ as a function of the convergence contributed by compact objects in the dark matter, $\kappa_{\text{b}}$, which may be primordial black holes or other dark matter compact objects but does not include the known contribution from intracluster stars.
@@ 5 Dependence on the microlens mass distribution
The upper limit derived in this paper on the contribution of compact objects to the dark matter actually extends over a very broad mass range. A remarkable result of the microlensing effect of compact masses is that the width of the corrugated network of MiCLs is simply $r_{w}=\kappa_{\star}/g$, and is independent of the mass distribution of the microlenses. The mass of the microlenses affects only the maximum magnification and frequency of the MiCL crossing events.
Let us first consider the case of compact objects in the dark matter that are much more massive than stars, with Einstein radius and convergence $\theta_{b}$ and $\kappa_{b}$. These compact objects should produce their own network of MiCLs, and the intracluster stars should produce a second level of corrugation on a smaller scale, corresponding to their smaller Einstein radius. Without the intracluster stars, the peak magnification of MiCL crossings within the network is typically (equations 20 and 27 of Venumadhav et al., 2017)
@@
Hence, our limit implies that no more than $\sim$ 3% of the dark matter can be in any compact objects above stellar masses, up to a maximum mass at which $\kappa_{b}=\kappa_{cb}=(g\theta_{b})^{2/3}$. For the upper limit $\kappa_{b}<0.02$, and the median value of the eigenvalue gradient for our cluster lenses $g^{-1}\simeq 16\,{\rm arcsec}$, this corresponds to $\theta_{b}\sim 0.05\,{\rm arcsec}$, or masses of $10^{9}$ to $10^{10}\,M_{\odot}$. Above this mass, the dark matter compact objects with $\kappa_{b}=0.02$ would produce isolated perturbations to the MaCL that could be individually identified around each object, which would stand out as a measurable perturbation caused by a dark object that is not associated with any luminous galaxy. Therefore, they would also be detectable in any case through their noticeable effect on cluster lensing models, above a minimum abundance that would need to be studied in more detail but which is unlikely to be close to the $\kappa_{cb}$ threshold for corrugation.
@@
However, there is in this case a different limitation to the detection of the microlensing effect: for any fixed $\kappa_{b}$, in the limit of very small masses of dark matter compact objects, the case of smooth dark matter must obviously be recovered. If the crossings of MiCLs can be detected down to a photometric accuracy of a fraction $f$ of the total flux, then the peak magnification of equation (4) must exceed $f\mu(r)$, where the mean magnification at angular separation $r$ from the MaCL is $\mu(r)\simeq|1-\kappa_{0}|^{-1}/(rg|\sin\alpha|)$ (equation 10 in Venumadhav et al., 2017), and $\alpha$ is the angle between the MaCL and the principal axis of the vanishing eigenvalue. Applying this to the typical separation where most highly magnified images should be seen, $r\sim r_{w}=(\kappa_{b}+\kappa_{\star})/g$, we obtain the condition
@@
For the case $\kappa_{b}\sim\kappa_{\star}\sim 0.01$, and $|\sin\alpha|\simeq 1/2$ and $f\sim 1/2$, we find that $D_{\rm S}\theta_{b}/R\gtrsim 2.5$, and for a typical stellar size and source angular diameter distance $R/D_{\rm S}\sim 10^{-15}$, we find $\theta_{b}\gtrsim 5\times 10^{-10}\,{\rm arcsec}$, implying that masses down to $10^{-6}\,M_{\odot}$ are still ruled out to contribute a surface density greater than $\kappa_{b}\gtrsim 0.02$ because any highly magnified, variable images would be observed over a width larger than observed around the MaCL.
@@
as expected at the separation $\sim r_{w}$ from the MaCL.
@@ 6 Conclusions
The result presented in this paper implies that the observation of only 9 supermagnified stars in lensing clusters of galaxies places an upper limit of 3% on the fraction of dark matter in compact objects which is already better than upper limits obtained so far from several microlensing surveys that have monitored millions of stars in the Magellanic Clouds (Alcock et al., 2001; Tisserand et al., 2007; Wyrzykowski et al., 2011b). Models of primordial black holes making up the dark matter are now ruled out for masses $m_{b}\gtrsim 10^{-6}\,M_{\odot}$ for fractions higher than $\sim$ 3% at the 99% confidence level. Interestingly, this limit is sensitive to the dark matter that is present in the most massive clusters of galaxies in the Universe, and also to the intergalactic dark matter that must be present along the line of sight to the source that traverses the lensing cluster and contributes to the total lensing convergence.
As the observations of variability in lensing clusters improve in the future with further observations of JWST and other telescopes, we can look forward to an improvement in these limits, and also to new constraints on the mass function of the microlenses caused by intracluster stars, or perhaps other unknown objects. Other models such as the QCD axion, which predicts the presence of axion minihalos that can also have lensing effects despite their extended mass distribution (Dai & Miralda-Escudé, 2020), will likely be constrained as well by monitoring campaigns of the supermagnified stars discovered by Kelly et al. (2018)
@@ 8 Data Availability
The data and code used for the work presented in this article are available at https://doi.org/10.6084/m9.figshare.25438648.v1.

### v2

@@ 4 Results
1 a MACS J1149 Icarus (Kelly et al., 2018) 0.54 1.49 0.832 12.46 0.0080 ++ 0.13 25-26 (F125W)
2 b MACS J0416 Spock-NW (Rodney et al., 2018) 0.397 1.0054 0.742 10.29 0.0028 25-26.5 (F814W)
@@
4 c MACS J0416 Warhol (Kaurov et al., 2019) 0.397 0.94 0.686 8.74 0.02 ++ 0.06 26.25 (F125W)
5 d WHL0137–08 Earendel (Welch et al., 2022) 0.566 6.2 0.590 24.01 0.0060 - 0.009 27.2 (F150W)
6 e PSZ1 G311.65 Godzilla (Diego et al., 2022) 0.443 2.37 ++ 0.55 $\approx$ 22 (F814W)
7 f Abell 2744 LS7 (Chen et al., 2022) 0.308 2.65 0.726 15.68 0.0052 - 0.15 27.05 (F150W)
8 g MACS J0647 star-1 (Meena et al., 2023b) 0.591 4.8 0.544 29.28 - 0.1 27.343 (F227W)
9 g MACS J0647 star-2 (Meena et al., 2023b) 0.591 4.8 0.549 30.59 - 0.32 28.330 (F227W)
10 h El Gordo Quyllur (Diego et al., 2023b) 0.870 2.1878 0.662 23.84 ++ 7.5 $\cdot 10^{-3}$ 25.5 (F356W)
11 i Abell 370 LS11 (Meena et al., 2023a) 0.375 1.2567 0.703 18.89 0.0017 - 0.6 29.51 (F200LP)
12 j MACS J0416 Mothra (Diego et al., 2023a) 0.397 2.091 0.936 15.78 - 0.05 $\approx$ 27.8 (F200W)
Table 1: List of 12 supermagnified star candidates, with lens and source redshifts $z_{L}$ and $z_{S}$, convergence at image position $\kappa_{0}$, eigenvalue inverse gradient $\mathbf{g}=\mathbf{\nabla}(-\kappa-\gamma$) at the macro critical curve (MaCL), angular separation from source to MaCL $r$ (with the positive direction pointing towards the outer side), and observed AB magnitude of each transient at given filter (AB mag). The values of $\kappa_{0}$ and $g^{-1}$ are the average of the set of publicly available lens models for each lens.
@@
    Caminha (v4) (Caminha et al. (2017)), GLAFIC (v4) (Oguri (2010), Kawamata et al. (2016), Kawamata et al. (2018)), Keeton (v4) (McCully et al. (2014), Ammons et al. (2014), Keeton (2010)), Sharon (v4 Cor.) (Johnson et al. (2014), Jullo et al. (2007)), Zitrin-ltm-gauss (v3) and Zitrin-nfw (v3) (Zitrin et al. (2009), Zitrin et al. (2013)).
@@
We now proceed to compute a maximum likelihood estimate of the optimal value of the total surface density of microlenses to explain the distribution of separations in Figure 1, and an upper limit to the fraction of dark matter that may be composed of compact objects. Initially, we assume for simplicity that any compact objects accounting for part of the dark matter have a similar mass distribution as the intracluster stars, and in the next section we discuss the range of masses that our limit actually applies to.
@@
where $p_{i}(r_{i},g_{i},\kappa_{\text{b}})$ is the probability density of observing a stellar microlensing event separated by $r_{i}$ from the MaCL of event $i$, given a microlensing band of width $r_{w,i}=(\kappa_{\star,i}+\kappa_{\text{b}})/g_{i}$. We assume that the total contribution from point masses to microlensing is the sum of the known intracluster light ($\kappa_{\star}$) and a hypothetical contribution from dark matter compact objects, such as primordial black holes ($\kappa_{\text{b}}$). We then find the optimal value of $\kappa_{\text{b}}$ that maximizes this likelihood function. We use all 9 events considered reliable. The value of $\kappa_{\star}$ is measured only for 5 of these events, and for the other 4 we simply use the average value of $\kappa_{\star}$ of all 7 events where it is measured (including events 2 and 3 in Table 1 that are not considered reliable as supermagnified stars), which is $\bar{\kappa}_{\star}=6\times 10^{-3}$.
@@
The result for the likelihood function is shown in Figure 2 as the solid, orange line. The optimal value of the dark matter surface density is consistent with zero, so the observed distribution of supermagnified star images is perfectly explained by the estimated stellar population surface density in the measured intracluster light, and is consistent with no compact objects in the dark matter. Figure 2 shows also the $(1,2,3)\sigma$ upper limits that can be placed on any additional compact objects, if we treat the $\ln\mathcal{L}$ function as a $\chi^{2}$ function. The $3\sigma$ upper limit, at a confidence level of 0.3%, is $\kappa_{b}<0.012$. We warn that this upper limit includes only the statistical uncertainty associated with the 9 events we used, but we have not incorporated other uncertainties due to the modeling of the lenses, the determination of $\kappa_{\star}$ or the assumed mass function of the stars in the intracluster light. This upper limit to the fraction of dark matter in compact objects should also be slightly increased when we take into account that $\sim$ 10% of the total mass contributing to lensing is in the hot, X-ray emitting gas (note that the gas usually has a density profile that is less centrally concentrated than the dark matter in the strong lensing region, so this fraction should be slightly smaller than the cosmic ratio of baryonic to total matter).
We show also as a blue line the result of ignoring the theoretical prediction that the corrugated band is centered on the SCL rather than the MaCL, and using equation (5) with $s$ directly replaced by $r$, rather than equation (6). In this case the upper limits would be increased.
@@
As discussed at the beginning of Section 3, we have excluded from this analysis event 6, which similarly to event 11 has a magnified image further from the model MaCL by a factor $\sim 4$ than the predicted width $r_{w}$ for $\kappa_{\star}=0.01\kappa_{0}$, because of the model uncertainty in the true separation from the MaCL. Had we included this event, the $3\sigma$ upper limit on $\kappa_{b}$ increases only to $\kappa_{b}<0.013$.
@@
As a consequence, we can establish an upper limit to the fraction of dark matter that may be in compact objects. The average local convergence in the 9 events we have used is $\bar{\kappa}_{0}=0.691$, so the $3\sigma$ upper limit on the dark matter fraction in compact objects is $\kappa_{b}/\bar{\kappa}_{0}\simeq 0.017$. If we take into account the fraction of the lens surface density $\kappa_{0}$ that is in the hot X-ray gas, this upper limit is only slightly increased to $\simeq 2\%$.
@@
Figure 2: Likelihood function $\ln\mathcal{L}$ as a function of the convergence contributed by compact objects in the dark matter, $\kappa_{\text{b}}$, which may be primordial black holes or other dark matter compact objects but does not include the known contribution from intracluster stars. The blue dashed line is obtained if we assume that the corrugated band is symmetric around the MaCL, while the orange line takes into account that the band is actually symmetric around the SCL.
@@ 5 Dependence on the microlens mass distribution
Although the upper limit on the contribution of compact objects by the dark matter has been derived assuming they have the same mass as intracluster stars, the result actually extends over a very broad mass range. A remarkable result of the microlensing effect of compact masses is that the width of the corrugated network of MiCLs is simply $r_{w}=\kappa_{\star}/g$, and is independent of the mass distribution of the microlenses. The mass of the microlenses affects only the maximum magnification and frequency of the MiCL crossing events.
To see this, let us first consider the case of compact objects in the dark matter that are much more massive than stars, with Einstein radius $\theta_{b}\gg\theta_{\star}$ and convergence $\kappa_{b}$. These compact objects should produce their own network of MiCLs, and the intracluster stars give rise to a second level of corrugation at a smaller scale of the microcaustics, corresponding to their smaller Einstein radius. Without the intracluster stars, the peak magnification of MiCL crossings within the network is typically (equations 20 and 27 of Venumadhav et al., 2017)
@@
Hence, our limit implies that no more than $\sim$ 2% of the dark matter can be in any compact objects above stellar masses, up to a maximum mass at which $\kappa_{b}=\kappa_{cb}=(g\theta_{b})^{2/3}$. For the upper limit $\kappa_{b}<0.012$, and the median value of the eigenvalue gradient for our cluster lenses $g^{-1}\simeq 16\,{\rm arcsec}$, this corresponds to $\theta_{b}\sim 0.05\,{\rm arcsec}$, or masses of $10^{9}$ to $10^{10}\,M_{\odot}$. Above this mass, the dark matter compact objects with $\kappa_{b}=0.012$ would produce isolated perturbations to the MaCL that could be individually identified around each object, which would stand out as a measurable perturbation caused by a dark object that is not associated with any luminous galaxy. Therefore, they would also be detectable in any case through their noticeable effect on cluster lensing models, above a minimum abundance that would depend on the specific observation and modeling but which is unlikely to be as high as the $\kappa_{cb}$ threshold for corrugation.
@@
However, there is in this case a different limitation to the detection of the microlensing effect: for any fixed $\kappa_{b}$, in the limit of very small masses of dark matter compact objects, the case of smooth dark matter must obviously be recovered. This occurs because the MiCL crossings overlap, as the separation between microcaustics is reduced below the radius of the source star. If the crossings of MiCLs can be detected down to a photometric accuracy of a fraction $f$ of the total flux, then the peak magnification of equation (8) must exceed $f\mu(r)$, where the mean magnification at angular separation $r$ from the MaCL is $\mu(r)\simeq|1-\kappa_{0}|^{-1}/(rg|\sin\alpha|)$ (equation 10 in Venumadhav et al., 2017), and $\alpha$ is the angle between the MaCL and the principal axis of the vanishing eigenvalue. Applying this to the typical separation where most highly magnified images should be seen, $r\sim r_{w}=(\kappa_{b}+\kappa_{\star})/g$, we obtain the condition
@@
For the case $\kappa_{b}\sim\kappa_{\star}\sim 0.01$, and $|\sin\alpha|\simeq 1/2$ and $f\sim 1/2$, we find that $D_{\rm S}\theta_{b}/R\gtrsim 2.5$, and for a typical stellar size and source angular diameter distance $R/D_{\rm S}\sim 10^{-15}$, we find $\theta_{b}\gtrsim 5\times 10^{-10}\,{\rm arcsec}$, implying that masses down to $10^{-6}\,M_{\odot}$ are still ruled out to contribute a surface density greater than $\kappa_{b}\gtrsim 0.012$ because any highly magnified, variable images would be observed over a width larger than observed around the MaCL.
@@
as expected at the separation $\sim r_{w}$ from the MaCL. For the minimum mass corresponding to the Einstein radius $\theta_{b}=(R/D_{s})\kappa_{b}^{3/2}(\kappa_{\star}+\kappa_{b})^{-2}|\sin\alpha|^{-2}$, the microcaustic peaks are overlapping over the whole corrugated network; in other words, the mean separation between neighboring microcaustics is equal to the source radius.
@@ 6 Conclusions
The result presented in this paper implies that the observation of only 9 supermagnified stars in lensing clusters of galaxies places an upper limit of 2% on the fraction of dark matter in compact objects. This is already better than upper limits obtained from several microlensing surveys that have monitored millions of stars in the Magellanic Clouds (Alcock et al., 2001; Tisserand et al., 2007; Wyrzykowski et al., 2011b), except for the recent OGLE result which achieves an even stronger upper limit (Mróz et al., 2024a). Models of primordial black holes making up the dark matter are now ruled out for masses $m_{b}\gtrsim 10^{-6}\,M_{\odot}$ for fractions higher than $\sim$ 2% at the 99% confidence level. Interestingly, this limit applies to the dark matter that is present in the most massive clusters of galaxies in the Universe, and also to the intergalactic dark matter that must be present along the line of sight to the source that traverses the lensing cluster, and contributes to the total lensing convergence.
As the observations of variability in lensing clusters improve in the future with further observations of JWST and other telescopes, we can look forward to an improvement in these limits, and also to new constraints on the mass function of the microlenses caused by intracluster stars, or perhaps other unknown objects. Other models such as the QCD axion, which predicts the presence of axion minihalos that can also have lensing effects despite their extended mass distribution (Dai & Miralda-Escudé, 2020), will likely be constrained as well by monitoring campaigns of the supermagnified stars discovered by Kelly et al. (2018).
@@ 8 Data Availability
The data and code used for the work presented in this article are available at https://doi.org/10.6084/m9.figshare.27635736.v1.